The most recent financial statements for Wise Co. are shown here:
Income Statement Balance Sheet
  Sales $ 53,200 Current assets $ 25,000 Long-term debt $ 53,500
  Costs 42,600 Fixed assets 96,000 Equity 67,500
  Taxable income $ 10,600   Total $ 121,000   Total $ 121,000
  Taxes (34%) 3,604
        
   Net income $ 6,996
    

Assets and costs are proportional to sales. The company maintains a constant 30 percent dividend payout ratio and a constant debt–equity ratio.

What is the maximum increase in sales that can be sustained (sustainable growth rate)? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  Maximum increase in sales $
Answer #1

ROE = Net Income / Total Equity = $6,996 / $67,500 = 0.1036, or 10.36%

Plowback Ratio = 1 - payout ratio

= 1 - 0.30 = 0.70

Sustainable Growth Rate = [ROE * b] / [1 - (ROE * b)]

= [0.1036 * 0.70] / [1 - (0.1036% * 0.70)]

= 0.0726 / 0.9274 = 0.0782, or 7.82%

Maximum Increase in Sales = Current Sales * g

= $53,200 * 0.0782 = $4,161.65